# A Simple Algorithm for Trimmed Multipoint Evaluation

### v1

@@ 1 Introduction
For univariate polynomials, a textbook algorithm [10, 26] solves the multipoint evaluation problem in near-linear time. This algorithm generalizes to multivariate polynomials [20] (via a simple divide-and-conquer method sometimes referred to as Yates’ algorithm [28]), however, only in the restricted setting where the evaluation points form a cartesian grid. Specifically, Yates’ algorithm evaluates an $n$-variate polynomials with individual degree $d$ on all points of a grid
@@
in near-optimal time [^1] Here and throughout, we write $O^{*}(\cdot)$ to omit polynomial factors in $n$ and $d$. $O^{*}((d+1)^{n})$. Lifting this restriction to grid points has been the focus of a long and active line of research [18, 23, 15, 25, 6, 24, 2, 1] which, following breakthroughs by Umans [23] and Kedlaya and Umans [15], only recently culminated in an algorithm with almost-optimal running time $(d+1)^{(1+o(1))n}\operatorname{poly}(n,d,\log|\mathbb{F}|)$, for all finite fields $\mathbb{F}$ and for $(d+1)^{n}$ arbitrary evaluation points, due to Bhargava, Ghosh, Guo, Kumar and Umans [1]. This fully settles the multipoint evaluation problem for dense polynomials (over finite fields), but leaves open whether almost-linear time can also be achieved for (some classes of) sparse polynomials and evaluation points.
In this paper we focus on one natural such class, called trimmed multipoint evaluation, with important applications in the design of exact and parameterized algorithms. Trimmed multipoint evaluation can be solved in near-linear time by an algorithm due to van der Hoeven and Schost [25]. Our contribution is that we make this result accessible to modern algorithm design (beyond computer algebra) by distilling a particularly simple recursive algorithm.
@@
to which we will informally refer as trimmed grids. This is arguably the most naturally matching class of evaluation points. To see this, first observe that the number of relevant (i.e., possibly nonzero) coefficients of $P$ equals exactly the number of grid points. We denote this number by $\binom{n}{\leq D}_{d}$ (which can be seen as an appropriate generalization of a binomial coefficient called an extended binomial coefficient). More importantly, it turns out that the polynomial $P$ is uniquely determined by the evaluations on an appropriate trimmed grid, i.e., we can interpolate $P$ given only these evaluations. Trimmed grids are, in a sense, the only sets of evaluation points satisfying this property; see [22].
@@
Besides being a natural problem in its own right, our interest in trimmed multipoint evaluation stems mainly from its applications in the context of exponential-time algorithms, most notably in a line of research on solving systems of polynomial equations [17, 7, 13, 12]. In this problem the input consists of $n$-variate polynomials $P_{1},\dots,P_{m}$ over some finite field $\mathbb{F}_{q}$ with (total) degree $\Delta$, and the task is to test if all polynomials simultaneously vanish at some point $x\in\mathbb{F}_{q}^{n}$, i.e., $P_{1}(x)=\dots=P_{m}(x)=0$. The initial breakthrough due to Lokshtanov, Paturi, Tamaki, Williams and Yu [17] established that this problem can be solved exponentially faster than brute-force, in time $(q-\epsilon)^{n}$ for some $\epsilon>0$, whenever $q$ and $\Delta$ are constant. This inspired several follow-up papers aiming to optimize the precise exponential running time [7, 13, 12]. Trimmed multipoint evaluation shows up as a critical subroutine in the algorithms due to Dinur [13] (for multilinear polynomials, $d=q-1=1$) and due to Dell, Haak, Kallmayer and Wennmann [12] (for general $d=q-1\geq 1$). Solving systems of polynomial equations in turn has many more applications on both the theoretical side—e.g., parity-counting directed Hamiltonian cycles [4]—and the practical side—e.g., the security of several so-called multivariate cryptosystems is based on the hardness of solving quadratic equations [13, 21, 16]. Especially for the latter it could be interesting to achieve simple, implementable algorithms.
In another closely related work, Björklund, Husfeldt, Kaski and Koivisto [5] considered trimmed [^3] In fact, they consider a more general definition of “trimmed” allowing arbitrary downward-closed sets. variants of the Zeta and Möbius transforms to develop fast algorithms for various exponential-time graph problems, such as computing the chromatic number for constant-degree-bounded graphs. The Zeta and Möbius transforms can be regarded as special cases of polynomial multipoint evaluation. [^4] Indeed, recall that the zeta transform of a function $f:\{0,1\}^{n}\to\mathbb{F}$ is defined as $(f\zeta)(X)=\sum_{Y\subseteq X}f(Y)$, where we identify sets $X\subseteq[n]$ with their indicator vectors $X\in\{0,1\}^{n}$. Consider the $n$-variate polynomial $P(x_{1},\dots,x_{n})=\sum_{Y\subseteq[n]}f(Y)\prod_{i\in Y}x_{i}$, and observe that $(f\zeta)(X)=P(X)$. Thus, the zeta transform $f\zeta$ can be read off the evaluations of $P$ on the grid $\{0,1\}^{n}$.
@@
The state of the art for trimmed multipoint evaluation, as mentioned before, is a clever algorithm with near-linear running time $O^{*}(\binom{n}{\leq D}_{d})$ due to van der Hoeven and Schost [25]. Their algorithm is based on the classical concept of Newton interpolation (see e.g. [3]), suitably tailored to the trimmed problem (see also [27, 9, 22, 11, 14] for some more references with a more mathematical point of view). The algorithm also offers two additional benefits: (1) It even solves a strictly more general multipoint evaluation problem on arbitrary downward-closed sets of relevant coefficients and grid points. (2) van der Hoeven and Schost have optimized the lower-order factors achieving an algebraic algorithm with $O(nN\log^{2}N\log\log N)$ field operations where $N=\binom{n}{\leq D}_{d}$. The potentially remaining lower-order improvements even persist in the univariate setting. Thus, all in all, the trimmed multipoint evaluation problem has already been satisfyingly resolved.
The only downside is that van der Hoeven and Schost’s algorithm is arguably somewhat intricate—both in the sense that it can be technically demanding to understand, particularly for researchers outside the computer algebra community, and in that it relies on several textbook algebraic primitives, such as efficient conversions between polynomial bases [3] and the use of truncated Fourier transforms as an implementation detail to achieve further improvements.
@@
Our focus here is not to optimize the running time or generality of this state of the art. Instead, given the many exciting algorithmic applications our contribution is to make van der Hoeven and Schost’s result for trimmed multipoint evaluation accessible to the algorithms community. We design a simple recursive algorithm that is teachable to researchers without any background in computer algebra. Moreover, our algorithm does not rely on any black-box algebraic primitives other than Gaussian elimination. As in [25], we also obtain an equally simple algorithm for the interpolation problem.
@@
The term “trimmed” multipoint evaluation is inspired by the related algorithms for Zeta and Möbius transforms in [5]. It is not standard in the mathematical literature, where one more often encounters terms like “triangular subsets of tensor product grids”. We adopt the “trimmed” terminology here as this paper is primarily intended for the algorithms community.
@@ 2 Preliminaries
We write $[n]=\{1,\dots,n\}$. Throughout, let $\mathbb{F}$ be a field and assume that we can evaluate field operations in unit time. For integers $n,k,d$ with $n\geq 0$ and $d\geq 1$ we define the extended binomial coefficient
@@
As a shorthand, we write $\binom{n}{\leq k}_{d}=\sum_{i=0}^{k}\binom{n}{i}_{d}$, and similarly define $\binom{[n]}{\leq k}_{d}=\bigcup_{i=0}^{k}\binom{n}{i}_{d}$. Further, we rely on the following generalization of Pascal’s triangle; see [8].
@@ 3.1 Key Ideas
Let $a\in\mathbb{F}^{d+1}$ be the coefficient vector of the univariate polynomial $P(X)=\sum_{i=0}^{d}a_{i}X^{i}$ of degree $d$. The evaluation of $P$ at all points $z_{j}$ with $j\in\{0,\dotsc,d\}$ can be expressed as the matrix-vector product $V\cdot a=y$ (indeed, each entry is of the form $y_{j}=\sum_{i=0}^{d}a_{i}z_{j}^{i}=P(z_{j})$). Similarly, we can interpolate $P$ from its evaluations $P(z_{j})$ using the matrix-vector product $V^{-1}\cdot y=V^{-1}\cdot V\cdot a=a$. The generalization to multivariate polynomials is simple: Let $a\in\mathbb{F}^{(d+1)^{n}}$ be the coefficient vector of an $n$-variate polynomial $P$ with individual degree $d$ that we want to evaluate on all grid points in $Z$. Then the Kronecker product $V(z_{1,0},\dotsc,z_{1,d})\otimes\cdots\otimes V(z_{n,0},\dotsc,z_{n,d})\cdot a=y$ yields the vector $y\in\mathbb{F}^{(d+1)^{n}}$ of all grid point evaluations. The classical algorithm of Yates [28] allows to compute these evaluations recursively. Notably, this computation does not account for the total degree $D$ and automatically results in a running time of $(d+1)^{n}$ for both multipoint evaluation and interpolation—even in our setting where the total degree $D$ is much smaller than its maximal value $nd$. Consequently, we need more insights to tailor this approach to the trimmed requirements.
@@
Our goal is to achieve a running time that is linear in $\binom{n}{\leq D}_{d}$, i.e., in the number of grid points on which we want to evaluate a multivariate polynomial $P$. As a baseline, we start with a short explanation of Yates’ algorithm: write $P(X_{1},\dots,X_{n})=\sum_{i=0}^{d}P_{i}(X_{1},\dots,X_{n-1})\cdot X_{n}^{i}$ to obtain $d+1$ many $(n-1)$-variate polynomials $P_{i}$ of degree $D-i$, where each $P_{i}$ can be seen as a coefficient of a univariate polynomial in $X_{n}$. By making $d+1$ many recursive calls of size $\binom{n-1}{\leq D}_{d}$, we compute the evaluations $P_{i}(Z_{\ell^{\prime}})$ for all $\ell^{\prime}\in\binom{[n-1]}{\leq D}_{d}$. We obtain all $P(Z_{\ell})$ for $\ell\in\binom{[n]}{\leq D}_{d}$ by evaluating the univariate polynomials $\sum_{i=0}^{d}P_{i}(Z_{\ell^{\prime}})\cdot X_{n}^{i}$ at all grid points $z_{n,i}$. However, this does not exploit the degrees $D-i$ of the polynomials $P_{i}$ in the recursive calls, and hence results in the running time of $(d+1)^{n}$.
Consider the following identity of the extended binomial coefficient that can be derived from Lemma 2.1
@@
In light of this identity, in order to achieve a running time of $O^{*}\big{(}\binom{n}{\leq D}_{d}\big{)}$ we aim to design an algorithm with $d+1$ recursive calls on $n-1$ variables with total degrees $D,D-1,\dotsc,D-d$.
@@
10       Recursively call TrimmedEvaluation$(Q_{j})$ to evaluate $Q_{j}(Z_{\ell^{\prime}})$ for $\ell^{\prime}\in\binom{[n-1]}{\leq D-j}_{d}$.
@@ 13       Let $k=\min\{d,D-\sum_{j=1}^{n-1}\ell^{\prime}_{j}\}$ and compute all evaluations
$\hskip 34.14322pt\begin{pmatrix}P\left\lparen Z_{(\ell^{\prime},0)}\right\rparen\\ \vdots\\ P\left\lparen Z_{(\ell^{\prime},k)}\right\rparen\\ \end{pmatrix}=\begin{pmatrix}L_{0,0}&&\\ \vdots&\ddots&\\ L_{k,0}&\cdots&L_{k,k}\\ \end{pmatrix}\cdot\begin{pmatrix}Q_{0}(Z_{\ell^{\prime}})\\ \vdots\\ Q_{k}(Z_{\ell^{\prime}})\\ \end{pmatrix}.$ ) .
Algorithm 1 TrimmedEvaluation$(P)$
@@
Our goal is an interleaving of the $P_{i}$’s into polynomials $Q_{j}$ satisfying two properties: (1) All evaluations $Q_{j}(Z_{\ell^{\prime}})$ can be recursively computed in a call of size $\binom{n-1}{\leq D-j}_{d}$; in particular, $Q_{j}$ must have degree $D-j$. (2) Simultaneously, we need to recover all evaluations $P(Z_{\ell})$ from the recursively computed evaluations $Q_{j}(Z_{\ell^{\prime}})$ for $\ell^{\prime}\in\binom{[n-1]}{\leq D-j}_{d}$. Due to the degree restriction for the polynomials $Q_{j}$ in (1) and evaluations $Q_{j}(Z_{\ell^{\prime}})$ in (2), this coincides with a matrix factorization of $V=L\cdot U$ where $U$ is upper-triangular and $L$ lower-triangular, i.e., a LU decomposition.
In more detail: Why is the LU decomposition useful in keeping the degrees of the polynomials $Q_{j}$ “small”? Let $p=(P_{0},\dots,P_{d})^{T}$ be the vector of polynomials $P_{i}$ and consider the product $U\cdot p=(Q_{1},\dotsc,Q_{d})^{T}=q$. Here, the upper triangular structure of $U$ guarantees that each $Q_{j}=\sum_{i=j}^{d}U_{j,i}\cdot P_{i}$ has at most degree $D-j$, since the first non-zero entry in the $j$-th row of $U$ corresponds to polynomial $P_{j}$ of degree $D-j$ while all other $P_{i}$’s with (possible) non-zero coefficients have a smaller degree. Additionally, we can also recover $P(Z_{\ell})$ from only the recursively computed $Q_{j}(Z_{\ell^{\prime}})$ as $L$ has lower triangular structure—appropriately exploiting the fact that $L\cdot q=L\cdot U\cdot p=V\cdot p$ [^5] The analogy to van der Hoeven and Schost’s algorithm is as follows: Broadly speaking, their algorithm is based on efficient transformations between not only the coefficient and evaluation-based representations of polynomials, but also involving a third representation based on so-called Newton polynomials. In this terminology, the matrix $U$ performs a basis change from the monomial basis to the Newton basis, and the matrix $L$ performs a basic change from the Newton basis to the evaluation basis. With these ideas in mind, we are now in the position to present the algorithm in detail.

### v2

@@ Abstract
keywords:
Algebraic Algorithms, Multipoint Evaluation, Interpolation, LU Decomposition
@@ 1 Introduction
For univariate polynomials, a textbook algorithm [BorodinM74, vonzurGathenG13] solves the multipoint evaluation problem in near-linear time. This algorithm generalizes to multivariate polynomials [Pan94] (via a simple divide-and-conquer method sometimes referred to as Yates’ algorithm [Yates37]), however, only in the restricted setting where the evaluation points form a cartesian grid. Specifically, Yates’ algorithm evaluates an $n$-variate polynomials with individual degree $d$ on all points of a grid
@@
in near-optimal time [^1] Here and throughout, we write $O^{*}(\cdot)$ to omit polynomial factors in $n$ and $d$. $O^{*}((d+1)^{n})$. Lifting this restriction to grid points has been the focus of a long and active line of research [NuskenZ04, Umans08, KedlayaU11, HoevenS13, BjorklundKW19, HoevenL21a, BhargavaG0M22, BhargavaGG0U22] which, following breakthroughs by Umans [Umans08] and Kedlaya and Umans [KedlayaU11], only recently culminated in an algorithm with almost-optimal running time $(d+1)^{(1+o(1))n}\operatorname{poly}(n,d,\log|\mathbb{F}|)$, for all finite fields $\mathbb{F}$ and for $(d+1)^{n}$ arbitrary evaluation points, due to Bhargava, Ghosh, Guo, Kumar and Umans [BhargavaGG0U22]. This fully settles the multipoint evaluation problem for dense polynomials (over finite fields), but leaves open whether almost-linear time can also be achieved for (some classes of) sparse polynomials and evaluation points.
In this paper we focus on one natural such class, called trimmed multipoint evaluation, with important applications in the design of exact and parameterized algorithms. Trimmed multipoint evaluation can be solved in near-linear time by an algorithm due to van der Hoeven and Schost [HoevenS13]. Our contribution is that we make this result accessible to modern algorithm design (beyond computer algebra) by distilling a particularly simple recursive algorithm.
@@
to which we will informally refer as trimmed grids. This is arguably the most naturally matching class of evaluation points. To see this, first observe that the number of relevant (i.e., possibly nonzero) coefficients of $P$ equals exactly the number of grid points. We denote this number by $\binom{n}{\leq D}_{d}$ (which can be seen as an appropriate generalization of a binomial coefficient called an extended binomial coefficient). More importantly, it turns out that the polynomial $P$ is uniquely determined by the evaluations on an appropriate trimmed grid, i.e., we can interpolate $P$ given only these evaluations. Trimmed grids are, in a sense, the only sets of evaluation points satisfying this property; see [Sauer04].
@@
Besides being a natural problem in its own right, our interest in trimmed multipoint evaluation stems mainly from its applications in the context of exponential-time algorithms, most notably in a line of research on solving systems of polynomial equations [LokshtanovPTWY17, BjorklundK019, Dinur21, DellHKW25]. In this problem the input consists of $n$-variate polynomials $P_{1},\dots,P_{m}$ over some finite field $\mathbb{F}_{q}$ with (total) degree $\Delta$, and the task is to test if all polynomials simultaneously vanish at some point $x\in\mathbb{F}_{q}^{n}$, i.e., $P_{1}(x)=\dots=P_{m}(x)=0$. The initial breakthrough due to Lokshtanov, Paturi, Tamaki, Williams and Yu [LokshtanovPTWY17] established that this problem can be solved exponentially faster than brute-force, in time $(q-\epsilon)^{n}$ for some $\epsilon>0$, whenever $q$ and $\Delta$ are constant. This inspired several follow-up papers aiming to optimize the precise exponential running time [BjorklundK019, Dinur21, DellHKW25]. Trimmed multipoint evaluation shows up as a critical subroutine in the algorithms due to Dinur [Dinur21] (for multilinear polynomials, d=q1=1d=q-1=1) and due to Dell, Haak, Kallmayer and Wennmann [DellHKW25] (for general $d=q-1\geq 1$). Solving systems of polynomial equations in turn has many more applications on both the theoretical side—e.g., parity-counting directed Hamiltonian cycles [BjorklundH13]—and the practical side—e.g., the security of several so-called multivariate cryptosystems is based on the hardness of solving quadratic equations [Dinur21, Patarin96, KipnisPG99]. Especially for the latter it could be interesting to achieve simple, implementable algorithms.
In another closely related work, Björklund, Husfeldt, Kaski and Koivisto [BjorklundHKK10] considered trimmed [^3] In fact, they consider a more general definition of “trimmed” allowing arbitrary downward-closed sets. variants of the Zeta and Möbius transforms to develop fast algorithms for various exponential-time graph problems, such as computing the chromatic number for constant-degree-bounded graphs. The Zeta and Möbius transforms can be regarded as special cases of polynomial multipoint evaluation. [^4] Indeed, recall that the zeta transform of a function $f:\{0,1\}^{n}\to\mathbb{F}$ is defined as $(f\zeta)(X)=\sum_{Y\subseteq X}f(Y)$, where we identify sets $X\subseteq[n]$ with their indicator vectors $X\in\{0,1\}^{n}$. Consider the $n$-variate polynomial $P(x_{1},\dots,x_{n})=\sum_{Y\subseteq[n]}f(Y)\prod_{i\in Y}x_{i}$, and observe that $(f\zeta)(X)=P(X)$. Thus, the zeta transform $f\zeta$ can be read off the evaluations of $P$ on the grid $\{0,1\}^{n}$.
@@
The state of the art for trimmed multipoint evaluation, as mentioned before, is a clever algorithm with near-linear running time $O^{*}(\binom{n}{\leq D}_{d})$ due to van der Hoeven and Schost [HoevenS13]. Their algorithm is based on the classical concept of Newton interpolation (see e.g. [BiniP94]), suitably tailored to the trimmed problem (see also [Werner80, Boor92, Sauer04, ChkifaCS14, DynF14] for some more references with a more mathematical point of view). The algorithm also offers two additional benefits: (1) It even solves a strictly more general multipoint evaluation problem on arbitrary downward-closed sets of relevant coefficients and grid points. (2) van der Hoeven and Schost have optimized the lower-order factors achieving an algebraic algorithm with $O(nN\log^{2}N\log\log N)$ field operations where $N=\binom{n}{\leq D}_{d}$. The potentially remaining lower-order improvements even persist in the univariate setting. Thus, all in all, the trimmed multipoint evaluation problem has already been satisfyingly resolved.
The only downside is that van der Hoeven and Schost’s algorithm is arguably somewhat intricate—both in the sense that it can be technically demanding to understand, particularly for researchers outside the computer algebra community, and in that it relies on several textbook algebraic primitives, such as efficient conversions between polynomial bases [BiniP94] and the use of truncated Fourier transforms as an implementation detail to achieve further improvements.
@@
Our focus here is not to optimize the running time or generality of this state of the art. Instead, given the many exciting algorithmic applications our contribution is to make van der Hoeven and Schost’s result for trimmed multipoint evaluation accessible to the algorithms community. We design a simple recursive algorithm that is teachable to researchers without any background in computer algebra. Moreover, our algorithm does not rely on any black-box algebraic primitives other than Gaussian elimination. As in [HoevenS13], we also obtain an equally simple algorithm for the interpolation problem.
@@
The term “trimmed” multipoint evaluation is inspired by the related algorithms for Zeta and Möbius transforms in [BjorklundHKK10]. It is not standard in the mathematical literature, where one more often encounters terms like “triangular subsets of tensor product grids”. We adopt the “trimmed” terminology here as this paper is primarily intended for the algorithms community.
@@ 2 Preliminaries
We write $n]=\{1,\dots,n\}$. Throughout, let $\mathbb{F}$ be a field and assume that we can evaluate field operations in unit time. For integers $n,k,d$ with $n\geq 0$ and $d\geq 1$ we define the extended binomial coefficient
@@
As a shorthand, we write $\binom{n}{\leq k}_{d}=\sum_{i=0}^{k}\binom{n}{i}_{d}$, and similarly define $\binom{[n]}{\leq k}_{d}=\bigcup_{i=0}^{k}\binom{n}{i}_{d}$. Further, we rely on the following generalization of Pascal’s triangle; see [Bollinger93].
@@ 3.1 Key Ideas
Let $a\in\mathbb{F}^{d+1}$ be the coefficient vector of the univariate polynomial $P(X)=\sum_{i=0}^{d}a_{i}X^{i}$ of degree $d$. The evaluation of $P$ at all points $z_{j}$ with $j\in\{0,\dotsc,d\}$ can be expressed as the matrix-vector product $V\cdot a=y$ (indeed, each entry is of the form $y_{j}=\sum_{i=0}^{d}a_{i}z_{j}^{i}=P(z_{j})$). Similarly, we can interpolate $P$ from its evaluations $P(z_{j})$ using the matrix-vector product $V^{-1}\cdot y=V^{-1}\cdot V\cdot a=a$. The generalization to multivariate polynomials is simple: Let $a\in\mathbb{F}^{(d+1)^{n}}$ be the coefficient vector of an $n$-variate polynomial $P$ with individual degree $d$ that we want to evaluate on all grid points in $Z$. Then the Kronecker product $V(z_{1,0},\dotsc,z_{1,d})\otimes\cdots\otimes V(z_{n,0},\dotsc,z_{n,d})\cdot a=y$ yields the vector $y\in\mathbb{F}^{(d+1)^{n}}$ of all grid point evaluations. The classical algorithm of Yates [Yates37] allows to compute these evaluations recursively. Notably, this computation does not account for the total degree $D$ and automatically results in a running time of $(d+1)^{n}$ for both multipoint evaluation and interpolation—even in our setting where the total degree $D$ is much smaller than its maximal value $nd$. Consequently, we need more insights to tailor this approach to the trimmed requirements.
@@
Our goal is to achieve a running time that is linear in $\binom{n}{\leq D}_{d}$, i.e., in the number of grid points on which we want to evaluate a multivariate polynomial $P$. As a baseline, we start with a short explanation of Yates’ algorithm: write $P(X_{1},\dots,X_{n})=\sum_{i=0}^{d}P_{i}(X_{1},\dots,X_{n-1})\cdot X_{n}^{i}$ to obtain $d+1$ many (n1)(n-1)-variate polynomials $P_{i}$ of degree DiD-i, where each $P_{i}$ can be seen as a coefficient of a univariate polynomial in $X_{n}$. By making $d+1$ many recursive calls of size $\binom{n-1}{\leq D}_{d}$, we compute the evaluations $P_{i}(Z_{\ell^{\prime}})$ for all $\ell^{\prime}\in\binom{[n-1]}{\leq D}_{d}$. We obtain all $P(Z_{\ell})$ for $\ell\in\binom{[n]}{\leq D}_{d}$ by evaluating the univariate polynomials $\sum_{i=0}^{d}P_{i}(Z_{\ell^{\prime}})\cdot X_{n}^{i}$ at all grid points $z_{n,i}$. However, this does not exploit the degrees DiD-i of the polynomials $P_{i}$ in the recursive calls, and hence results in the running time of $(d+1)^{n}$.
Consider the following identity of the extended binomial coefficient that can be derived from Lemma˜2.1
@@
In light of this identity, in order to achieve a running time of $O^{*}\big(\binom{n}{\leq D}_{d}\big)$ we aim to design an algorithm with $d+1$ recursive calls on n1n-1 variables with total degrees $D,D-1,\dotsc,D-d$.
@@
[…]
@@ 13   Let $k=\min\{d,D-\sum_{j=1}^{n-1}\ell^{\prime}_{j}\}$ and compute all evaluations
[…]
Algorithm 1 TrimmedEvaluation(P)(P)
@@
Our goal is an interleaving of the $P_{i}$’s into polynomials $Q_{j}$ satisfying two properties: (1) All evaluations $Q_{j}(Z_{\ell^{\prime}})$ can be recursively computed in a call of size $\binom{n-1}{\leq D-j}_{d}$; in particular, $Q_{j}$ must have degree DjD-j. (2) Simultaneously, we need to recover all evaluations $P(Z_{\ell})$ from the recursively computed evaluations $Q_{j}(Z_{\ell^{\prime}})$ for $\ell^{\prime}\in\binom{[n-1]}{\leq D-j}_{d}$. Due to the degree restriction for the polynomials $Q_{j}$ in (1) and evaluations $Q_{j}(Z_{\ell^{\prime}})$ in (2), this coincides with a matrix factorization of $V=L\cdot U$ where $U$ is upper-triangular and $L$ lower-triangular, i.e., a LU decomposition.
In more detail: Why is the LU decomposition useful in keeping the degrees of the polynomials $Q_{j}$ “small”? Let $p=(P_{0},\dots,P_{d})^{T}$ be the vector of polynomials $P_{i}$ and consider the product $U\cdot p=(Q_{1},\dotsc,Q_{d})^{T}=q$. Here, the upper triangular structure of $U$ guarantees that each $Q_{j}=\sum_{i=j}^{d}U_{j,i}\cdot P_{i}$ has at most degree DjD-j, since the first non-zero entry in the $j$-th row of $U$ corresponds to polynomial $P_{j}$ of degree DjD-j while all other $P_{i}$’s with (possible) non-zero coefficients have a smaller degree. Additionally, we can also recover $P(Z_{\ell})$ from only the recursively computed $Q_{j}(Z_{\ell^{\prime}})$ as $L$ has lower triangular structure—appropriately exploiting the fact that $L\cdot q=L\cdot U\cdot p=V\cdot p$. [^5] The analogy to van der Hoeven and Schost’s algorithm is as follows: Broadly speaking, their algorithm is based on efficient transformations between not only the coefficient and evaluation-based representations of polynomials, but also involving a third representation based on so-called Newton polynomials. In this terminology, the matrix $U$ performs a basis change from the monomial basis to the Newton basis, and the matrix $L$ performs a basic change from the Newton basis to the evaluation basis. With these ideas in mind, we are now in the position to present the algorithm in detail.